The square is divided into 5 congruent rectangles. If the perimeter of one rectangle is 30 units, what is the perimeter of the square?
Answer by RAY100(1637)

A rough sketch shows a square, of side L, with 5 rectangles inside, each L by L/5
.
Perimeter of rectangles = 2L +2(L/5) = 2.4 L = 30,,,,L= 12.5
.
Perimeter of square = 4*L=4*12.5 = 50 units
.